Express one ton of refrigeration in terms of kJ/hour.


1
Expert's answer
2021-02-15T05:40:49-0500

Q163204


Express one ton of refrigeration in terms of kJ/hour.


Solution:


One ton of Refrigeration: It is a unit of power used for describing the heat extracting capacity of refrigerators and Air conditioning equipment.


One ton Refrigeration is the amount of heat extracted per day for converting 1 ton (907 kg) of water at 0 0C to Ice at 0 0C.


The Heat of fusion of water is 334 J/g.

and 907 kg in grams "= 907 kg * \frac{1000g}{1kg} = 907,000 grams"


So for converting 907,000 g of water ( 0 0C) to ice ( 0 0 C), we will have to extract


Q = m * ΔH fusion = 907,000 g * 334 J/g = 302,938,000 J.


So we need to extract 302,938,000 Joules of heat in a day for converting 1 ton ( 907g ) of

water at 0 0C to ice at 0 0C.


1-ton Refrigeration = 302,938,000 J / day.


We have to convert this to kJ / hour.


We will use the conversion factor 1kJ = 1000J and 1 day = 24 hours for converting

this to kJ/hour


1-ton refrigeration in kJ/hour "= \frac{302,938,000\space Joule}{1\space day}*\frac{1\space kJ}{1000\space J} * \frac{1\space day}{24\space hours }= 12,622 kJ/hour."



So one-ton of refrigeration is equal to 12,622 kJ/hour.


(Note: In the USA 1 short ton (907 kg) is used for measuring one-ton refrigeration.

In some countries, 1 ton (1000 kg water ) is used for measuring one-ton refrigeration.

After plugging 1000 kg instead of 907 kg we will get 13,917 kJ/hr as the answer. )